# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v1

@@ 7 Empirical Validation of Predicted Depth and Width Scaling
Figure 2: Strong linear fits imply theory/experiment match for modeling the impact of depth (left, $d=4.8\log_{2}n-15.8$ with $r^{2}=0.93$) and width (right, $n=7.2\log_{2}w-41.7$ with $r^{2}=0.98$) on effective context length for the $A_{5}$ state tracking task, a canonical hard regular language recognition problem. As predicted by Thms. 4.1 and 6.1, to recognize strings of length $n$, depth only needs to increase minimally $\propto\log n$ while width must increase drastically as $\exp(\Theta(n))$.
@@
We report on an extensive set of experiments to address these questions, training models of different depths and widths on the $A_{5}$ state tracking task (Merrill et al., 2024), which is a canonical testbed for hard regular language recognition (Thm. 4.1). The input to the task is a sequence of elements in $A_{5}$ (the group of even permutations over 5 elements), and the label at each token is the cumulative product of previous permutations up to and including that token (which is itself an element of $A_{5}$).
@@
We train several (non-universal) transformers with the same architecture used by Merrill et al. (2024) on about 100 million sequences of the $A_{5}$ task of varying lengths up to 1024 (this took about 1000 GPU hours). To understand the impact of depth and width in a controlled way, we train two series of transformers: the first with width fixed to 512 and depth varying in $\{6,9,12,15,18,21,24\}$, and the second with depth fixed to 6 and width varying in $\{128,256,512,1024\}$. See Appendix E for further details about our training procedure. After each model is trained, we measure accuracy at each token index from 1 to 1024 and define $n^{*}$ as the maximum token index at which the model achieved at least 95% validation accuracy. As we trained several seeds with the same depth and width, we aggregate these results across all models with the same depth and width by taking the best-performing (max $n^{*}$) model. We then plot $n^{*}$, which represents the effective context length up to which a model can solve the $A_{5}$ problem, as a function of either depth or width, holding the other variable fixed. We then evaluate if the predicted theoretical relationships between depth, width, and context length hold via an $r^{2}$ statistic.
@@
The results are shown in Fig. 2. When varying depth (left plot), there is a very strong positive correlation ($r^{2}=0.93$) between model depth (x-axis) and $\log n^{*}$ (y-axis, log scale), the effective (log) context length till which it can solve problems with high accuracy. When varying width (right plot) there is an even stronger positive correlation ($r^{2}=0.98$) between log width (x-axis, log scale) and $n^{*}$ (y-axis). These results provide strong empirical support for our theoretical predictions that, to recognize regular languages over strings of length $n$, increasing depth logarithmically in $n$ will suffice (Thm. 4.1), but width must increase exponentially in $n$ (Thm. 6.1). Fig. 2 also gives us a strongly predictive functional form to quantify the impact of scaling depth or width on the effective context length for regular language recognition. The empirical slope for the depth relationship is is 4.8 layers per log tokens. This is less than the slope of 8 derived for universal transformers in Thm. 4.1, but slightly greater than the theoretical coefficient of 4 for transformers whose depth grows non-uniformly with context length. Thus, our transformers have learned a construction whose depth coefficient is comparable to what we showed was possible in theory, though perhaps slightly more wasteful than it needs to be. Overall, these empirical results show that, in practice, the impact of depth and width on effective context length for regular language recognition aligns with our theoretical predictions, and we are able to empirically fit the quantitative coefficients in the relationships.
@@ 8 Conclusion
We have shown that recognizing regular languages and graph connectivity, two key problems inexpressible by fixed-depth transformers, become expressible if the depth of the transformer can grow very slightly (logarithmically) with the context length. This implies transformers with fixed depth $d$ can solve these problems up to bounded context lengths of $2^{O(d)}$. Further, we showed that scaling depth to solve these problems is more efficient than scaling width (which requires superpolynomial increase) or scaling chain-of-thought steps (which requires superlogarithmic increase). As dynamic test-time compute methods have become popular for building more powerful reasoning models such as OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek AI, 2025), it would be interesting to explore whether universal transformers can realize this theoretical efficiency to provide more efficient long-context reasoning than chain-of-thought steps in practice.
Limitations of Log Depth.
While we have shown that growing depth logarithmically allows transformers to express some interesting problems that fixed-depth transformers could not, there are, in general, limitations on the types of problems that can be made expressible using log depth. We know that log-depth transformers can be simulated in $\mathsf{TC}^{1}$. Thus, unless $\mathsf{NC}=\mathsf{P}$, log-depth (or even polylog-depth, i.e., $\log^{k}n$) transformers cannot express $\mathsf{P}$-complete problems including solving linear equalities, in-context context-free language recognition (given both a grammar $G$ and string $x$ as input, does $G$ generate $x$?), circuit evaluation, and determining the satisfiability of Horn clauses (Greenlaw et al., 1991). Beyond $\mathsf{P}$-complete problems, other natural problems could be inexpressible by log-depth transformers. Interesting candidates include context-free recognition (generalizing regular languages; Thm. 4.1), which is in $\mathsf{NC}^{2}$ (Ruzzo, 1981) and even boolean formula evaluation, which is in $\mathsf{NC}^{1}$. In future work, it would be interesting to empirically and theoretically study the depth required for transformers to solve these problems.
Impact Statement
This paper presents work whose goal is to advance our foundational understanding of the transformer architecture widely used in modern deep learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
We appreciate Sanjeev Arora’s suggestion to consider the effect of scaling depth on transformers’ expressive power, rather than abstracting it away as a constant. We also benefited from insightful discussion during the Simons Institute workshop on Transformers as a Computational Model that motivated Thm. 6.2, Jackson Petty’s assistance with his state tracking codebase, and valuable input from anonymous reviewers. We also appreciate comments from Volodymyr Tkachuk. This project was supported in part by NSF award 1922658, and WM was supported by an NSF graduate research fellowship, Two Sigma PhD fellowship, and Ai2.
@@ A.1 Residual Stream Storage Interface
The first layer computes $z$ and stores $\psi(z)$ at blank indices $I$ in the residual stream, producing $\mathbf{h}^{\prime}$. Thus, the second layer can read $\phi(z)$ with masked pre-norm via Lem. A.1 and can also recompute $z$ from $\mathbf{h}$, which is a subspace of $\mathbf{h}^{\prime}$. At this point, it outputs $-\psi(z)+\phi(z)$ at indices $I$, which leads to $\mathbf{h}^{\prime\prime}$ storing $\phi(z)$ at $I$. ∎
@@ A.2 Clearing Stored Values
Combining Lem. A.4 with a parallel layer that stores some new value at $I$, we see that we can effectively overwrite values at $I$ rather than just deleting them.
@@ A.3 Computing Position Offsets
The precondition that we can identify the initial token (cf. Merrill & Sabharwal, 2024) is easy to meet with any natural representation of position, including $1/i$ or $\phi(i)$, as we can simply compare the position representation against some constant.
We assume that the positional encodings used by the model allow detecting the initial token (Merrill & Sabharwal, 2024). One way to enable this would simply be to add a beginning-of-sequence token, although most position embeddings should also enable it directly.
@@ Appendix B Division Construction Correctness
The proof of Lem. 3.1 presents the full construction to implement division in a transformer. For space, we omitted a full proof of correctness for the construction, which we now present.
@@
In the second layer, $q_{i}\cdot k_{j}=e_{j}-[\phi(j)]_{0}$ where $[\phi(j)]_{0}=j/\sqrt{2j^{2}+2}$ is the first coordinate of $\phi(j)$. Note that $[\phi(j)]_{0}\in[0,1)$ for positions $j\leq i$ and that it is monotonically increasing in $j$. It follows that the dot product is maximized at the largest $j\leq i$ such that $e_{j}=1$, i.e., at the largest $j\leq i$ that is a multiple of $m$. This $j$ has the property that $\lfloor i/m\rfloor=j/m$. Thus, the head at this layer attends solely to this $j$ and retrieves the value $\phi(j/m)=\phi(\lfloor i/m\rfloor)$ as intended.
@@
In the fifth layer, $q_{i}\cdot k_{j}\leq 2$ and the dot product achieves this upper limit exactly when two conditions hold: $b^{1}_{j}=1$ and $\lfloor i/m\rfloor=\lfloor j/m\rfloor$. Thus, as desired, the head at $i$ attends to all positions $j\leq i$ that have the same quotient as $i$ and also have $b^{1}_{j}=1$. Write $i$ as $i=b^{\prime}m+c^{\prime}$ for some $c^{\prime}<m$. It follows that the query-key dot product is maximized precisely at the $c^{\prime}$ positions $b^{\prime}m+1,b^{\prime}m+2,\ldots,b^{\prime}m+c^{\prime}$. Of these positions, only $b^{\prime}m+1$ has the property that the quotient there is not the same as the quotient two position earlier, as captured by the value $v_{j}=1-b^{2}_{j}$. Thus, the value $v_{j}$ is $1$ among these positions only at $j=b^{\prime}m+1$, and 0 elsewhere. The head thus attends uniformly at $c^{\prime}$ positions and retrieves $1/c^{\prime}$. By construction, $c^{\prime}=i-b^{\prime}m=i-\lfloor i/m\rfloor m$, showing that this layer also behaves as intended.
@@
Finally, that the sixth and seventh layers operate as desired is easy to see from the construction. ∎
Appendix C Graph Connectivity Proof
See 5.1
@@ Appendix C Graph Connectivity Proof
The idea is to use the first $n^{2}$ tokens of the transformer to construct binary predicates $B_{\ell}(i,j)$ for $\ell\in\{0,1,\ldots,\lceil\log n\rceil\}$ capturing whether $G$ has a path of length at most $2^{\ell}$ from $i$ to $j$. To this end, the transformer will use the $n^{3}$ padding tokens to also construct intermediate ternary predicates $C_{\ell}(i,k,j)$ for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$ capturing whether $G$ has paths of length at most $2^{\ell-1}$ from $i$ to $k$ and from $k$ to $j$. These two series of predicates are computed from each other iteratively:
@@
We next describe how the transformer operationalizes the computation of predicates $B_{\ell}$ and $C_{\ell}$. The input to the transformer is the adjacency matrix $A$ represented using $n^{2}$ tokens from $\{0,1\}$, followed by $n^{3}$ padding tokens $\square$, and finally the source and target nodes $s,t\in\{1,\ldots,n\}$ represented in unary notation using special tokens $a$ and $b$:
$A_{1,1}\ldots A_{1,n}\ A_{2,1}\ldots A_{2,n}\ \ldots\ldots\ A_{n,1}\ldots A_{n,n}$
$\underbrace{\square\ldots\ldots\ldots\square}_{n^{3}}\ \underbrace{a\ldots\ldots a}_{s}\ \underbrace{b\ldots\ldots b}_{t}$
@@
The transformer uses the next 15 layers to compute and store in the residual stream the semantic “coordinates” of each of the first $n^{2}+n^{3}$ token position as follows. For each of the first $n^{2}$ positions $p=in+j$ with $1\leq p\leq n^{2}$, it uses Lem. 3.1 (7 layers) with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi(i)$ and $\phi(j)$ to the residual stream at position $p$. In parallel, for each of the next $n^{3}$ positions $p=n^{2}+(in^{2}+kn+j)$ with $n^{2}+1\leq p\leq n^{2}+n^{3}$, it uses Lem. 3.1 with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi((i+1)n+k)$ and $\phi(j)$ to the residual stream. It then uses the lemma again (7 more layers), this time with $a_{i}$ set to $(i+1)n+k$ and $m$ again set to $n$, to add $\phi(i+1)$ and $\phi(k)$ to the residual stream. Lastly, it uses Lem. A.6 applied to $\phi(i+1)$ to add $\phi(i)$ to the residual stream.
Layer 17 of the transformer computes the predicate $B_{0}(i,j)$ at the first $n^{2}$ token positions as follows. At position $p=in+j$, it uses Lem. A.7 to compute $\mathbb{I}(\phi(A(i,j)=\phi(1))$ and $\mathbb{I}(\phi(i)=\phi(j))$; note that $\phi(A(i,j))$, $\phi(i)$, and $\phi(j)$ are available in the residual stream at position $p$. It then uses a feedforward layer to output $1$ if both of these are $1$, and output 00 otherwise. This is precisely the intended value of $B_{0}(i,j)$. The sublayer then adds $B_{0}(i,j)$ to the residual stream. The layer also adds to the residual stream the value $1$, which will be used to initialize the boolean that controls layer alternation in the repeated layers as discussed next.
@@
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $(2\ell-1)+8$ of the transformer computes the predicate $C_{\ell}(i,k,j)$ at the set of $n^{3}$ (padding) positions $p=n^{2}+in^{2}+kn+j$, as follows. It uses two heads, one with query $\langle\phi(i),\phi(k)\rangle$ and the other with query $\langle\phi(k),\phi(j)\rangle$. The keys in the first $n^{2}$ positions $q=i^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$, and the values are set to $B_{\ell-1}(i^{\prime},j^{\prime})$. The two heads thus attend solely to positions with coordinates $(i,k)$ and $(k,j)$, respectively, and retrieve boolean values $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$, respectively, stored there in the previous layer. The layer then uses Lem. A.7 to compute $\mathbb{I}(B_{\ell-1}(i,k)=1)$ and $\mathbb{I}(B_{\ell-1}(k,j)=1)$, and uses a feedforward layer to output $1$ if both of these checks pass, and output 00 otherwise. This is precisely the intended value of $C_{\ell}(i,k,j)$. If $\ell>1$, the layer replaces the value $C_{\ell-1}(i,k,j)$ stored previously in the residual stream with the new boolean value $C_{\ell}(i,k,j)$ by adding $C_{\ell}(i,k,j)-C_{\ell-1}(i,k,j)$ to the same coordinates of the residual stream. If $\ell=1$, it simply adds $C_{\ell}(i,k,j)$ to the residual stream.

### v2

@@ 7 Experiments: Learning to Recognize Regular Languages
Figure 2: Strong linear fits imply theory/experiment match for modeling the impact of depth (left, $d=4.8\log_{2}n-15.8$ with $r^{2}=0.93$) and width (right, $n=7.2\log_{2}w-41.7$ with $r^{2}=0.98$) on effective context length for the $A_{5}$ state tracking task, a canonical hard regular language recognition problem. As predicted by Theorems 1 and 3, to recognize strings of length $n$, depth only needs to increase minimally $\propto\log n$ while width must increase drastically as $\exp(\Theta(n))$.
@@
We report on an extensive set of experiments to address these questions, training models of different depths and widths on the $A_{5}$ state tracking task (Merrill et al., 2024), which is a canonical testbed for hard regular language recognition (Theorem 1). The input to the task is a sequence of elements in $A_{5}$ (the group of even permutations over 5 elements), and the label at each token is the cumulative product of previous permutations up to and including that token (which is itself an element of $A_{5}$).
@@
We train several (non-universal) transformers with the same architecture used by Merrill et al. (2024) on  100 million $A_{5}$ sequences of varying lengths up to 1024. To understand the impact of depth and width in a controlled way, we train two series of transformers: the first with width fixed to 512 and depth varying in $\{6,9,12,15,18,21,24\}$, and the second with depth fixed to 6 and width varying in $\{128,256,512,1024\}$. See Appendix F for further details about our training procedure. After each model is trained, we measure accuracy at each token index from 1 to 1024 and define $n^{*}$ as the maximum token index at which the model achieved at least 95% validation accuracy. As we trained several seeds with the same depth and width, we aggregate these results across all models with the same depth and width by taking the best-performing (max $n^{*}$) model. We then plot $n^{*}$, which represents the effective context length up to which a model can solve the $A_{5}$ problem, as a function of either depth or width, holding the other variable fixed. We then evaluate if the predicted theoretical relationships between depth, width, and context length hold via an $r^{2}$ statistic.
@@
The results are shown in Figure 2. When varying depth (left plot), there is a very strong positive correlation ($r^{2}=0.93$) between model depth (x-axis) and $\log n^{*}$ (y-axis, log scale), the effective (log) context length till which it can solve problems with high accuracy. When varying width (right plot) there is an even stronger positive correlation ($r^{2}=0.98$) between log width (x-axis, log scale) and $n^{*}$ (y-axis). These results provide strong empirical support for our theoretical predictions that, to recognize regular languages over strings of length $n$, increasing depth logarithmically in $n$ will suffice (Theorem 1), but width must increase exponentially in $n$ (Theorem 3). Figure 2 also gives us a strongly predictive functional form to quantify the impact of scaling depth or width on the effective context length for regular language recognition. The empirical slope for the depth relationship is is 4.8 layers per log tokens. This is less than the slope of 8 derived for universal transformers in Theorem 1, but slightly greater than the theoretical coefficient of 4 for transformers whose depth grows non-uniformly with context length. Thus, our transformers have learned a construction whose depth coefficient is comparable to what we showed was possible in theory, though perhaps slightly more wasteful than it needs to be. Overall, these empirical results show that, in practice, the impact of depth and width on effective context length for regular language recognition aligns with our theoretical predictions, and we are able to empirically fit the quantitative coefficients in the relationships.
@@ 8 Conclusion
We have shown that recognizing regular languages and graph connectivity, two key problems inexpressible by fixed-depth transformers, become expressible if the depth of the transformer can grow very slightly (logarithmically) with the context length by repeating layers. This implies transformers with fixed depth $d$ can solve these problems up to bounded context lengths of $2^{O(d)}$. Further, we showed that scaling depth to solve these problems is more efficient than scaling width (which requires superpolynomial increase) or scaling chain-of-thought steps (which requires superlogarithmic increase). As dynamic test-time compute methods have become popular for building more powerful reasoning models such as OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek AI, 2025), it would be interesting to explore whether universal transformers can realize this theoretical efficiency to provide more efficient long-context reasoning than chain-of-thought steps in practice.
While growing depth enables transformers to solve some key problems outside $\mathsf{TC}^{0}$, there are limitations on the types of problems log depth should enable solving. Unless $\mathsf{NC}=\mathsf{P}$, log-depth (or even polylog-depth) transformers cannot express $\mathsf{P}$-complete problems including solving linear equalities, in-context context-free language recognition, circuit evaluation, and determining the satisfiability of Horn clauses (Greenlaw et al., 1991). In future work, it would interesting to study the depth required for transformers to solve other interesting problems in $\mathsf{NC}$ including context-free recognition (generalizing regular languages; Theorem 1), which is in $\mathsf{AC}^{1}$ (Ruzzo, 1981; Venkateswaran, 1991) and boolean formula evaluation, which is $\mathsf{NC}^{1}$-complete (Buss, 1987). This would help us better understand the degree to which repeating layers can be used as a form of interence-time compute.
@@ A.1 Residual Stream Storage Interface
The first layer computes $z$ and stores $\psi(z)$ at blank indices $I$ in the residual stream, producing $\mathbf{h}^{\prime}$. Thus, the second layer can read $\phi(z)$ with masked pre-norm via Lemma 2 and can also recompute $z$ from $\mathbf{h}$, which is a subspace of $\mathbf{h}^{\prime}$. At this point, it outputs $-\psi(z)+\phi(z)$ at indices $I$, which leads to $\mathbf{h}^{\prime\prime}$ storing $\phi(z)$ at $I$. ∎
@@ A.2 Clearing Stored Values
Combining Lemma 5 with a parallel layer that stores some new value at $I$, we see that we can effectively overwrite values at $I$ rather than just deleting them.
@@ A.3 Computing Position Offsets
The precondition that we can identify the initial token [cf. Merrill and Sabharwal, 2024] is easy to meet with any natural representation of position, including $1/i$ or $\phi(i)$, as we can simply compare the position representation against some constant.
We assume that the positional encodings used by the model allow detecting the initial token [Merrill and Sabharwal, 2024]. One way to enable this would simply be to add a beginning-of-sequence token, although most position embeddings should also enable it directly.
@@ Appendix B Division Construction Correctness
The proof of Lemma 1 presents the full construction to implement division in a transformer. For space, we omitted a full proof of correctness for the construction, which we now present.
@@
For the second layer, let $h^{2}_{i}$ denote the head’s output at position $i$. By construction, $q_{i}\cdot k_{j}=e_{j}-[\phi(j)]_{0}$ where $[\phi(j)]_{0}=j/\sqrt{2j^{2}+2}$ is the first coordinate of $\phi(j)$. Note that $[\phi(j)]_{0}\in[0,1)$ for positions $j\leq i$ and that it is monotonically increasing in $j$. It follows that the dot product is maximized at the largest $j\leq i$ such that $e_{j}=1$, i.e., at the largest $j\leq i$ that is a multiple of $m$. This $j$ has the property that $\lfloor i/m\rfloor=j/m$. Thus, the head at this layer attends solely to this $j$. Recall that the value $v_{j}$ at this position is $h^{1}_{j}=\phi(j/m)$, which thus equals $\phi(\lfloor i/m\rfloor)$. The head’s output, $h^{2}_{i}$, is therefore $\phi(\lfloor i/m\rfloor)$, as intended, which is stored in the residual stream.
@@
In the fifth layer, position $i$ attends with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$; each $b^{k}_{j}$ for $k\in\{1,2\}$ can be retrieved from the residual stream. The query-key product achieves its upper bound of 2 exactly when two conditions hold: $\lfloor i/m\rfloor=\lfloor j/m\rfloor$ and $b^{1}_{j}=1$. Thus, the head attends from $i$ to all $j\leq i$ that store the same quotient as $i$ and also have $b^{1}_{j}=1$. To make this clearer, let’s write $i$ as $i=b^{\prime}m+c^{\prime}$ for some $c^{\prime}<m$. The query-key dot product is then maximized precisely at the $c^{\prime}$ positions $j$ in $\{b^{\prime}m+1,b^{\prime}m+2,\ldots,b^{\prime}m+c^{\prime}\}$, for all of which $\lfloor j/m\rfloor=\lfloor i/m\rfloor=b^{\prime}$; note that $b^{\prime}m$ is not included in this list as $b^{1}_{j}=0$ when $j=b^{\prime}m$. Of these positions, only $b^{\prime}m+1$ has the property that the quotient there is not the same as the quotient two position earlier, as captured by the value $v_{j}=1-b^{2}_{j}$. Thus, the value $v_{j}$ is $1$ among these positions only at $j=b^{\prime}m+1$, and 0 elsewhere.
Assuming $m$ does not divide $i$, $c^{\prime}>0$ and the head attends uniformly at $c^{\prime}$ positions, returning $1/c^{\prime}$ as the head output. By construction, $c^{\prime}=i-b^{\prime}m=i\bmod m$. The layer adds the vector $\psi(1,1/c^{\prime})$ defined as $\langle 1,1/c^{\prime},-1,-1/c^{\prime}\rangle$ to the residual stream at position $i$. This, when read in the next layer using masked pre-norm, will yield $\phi(1,1/c^{\prime})=\phi(c^{\prime})=\phi(i\bmod m)$.
On the other hand, if $m$ does divide $i$ (which can be checked with a separate, parallel head), we write $\psi(0)$ to the residual stream, which, when read by the next layer, will yield $\phi(0)=\phi(i\bmod m)$.
The sixth layer attends with query $q_{i}=\phi(a_{i})$, key $k_{j}=\phi(j)$, and value $v_{j}=\langle\phi(\lfloor j/m\rfloor),\phi(j\bmod m)\rangle$, where both components of the value have been previously computed and stored in the residual stream in layers two and five. Since $a_{i}\leq i$, the query matches the key at exactly one position $j=a_{i}$, and the head retrieves $\langle\phi(\lfloor a_{i}/m\rfloor),\phi(a_{i}\bmod m)\rangle$, which is precisely $\langle\phi(b_{i}),\phi(c_{i})\rangle$. The layer can thus store $\phi(b_{i})$ and $\phi(c_{i})$ to the residual stream at position $i$.
@@
That the seventh “cleanup” layer operates as desired is easy to see from the construction. ∎
Appendix C Regular Language Recognition Proof
See 1
Proof.
Regular language recognition can be framed as multiplying a sequence of elements in the automaton’s transition monoid [Myhill, 1957, Thérien, 1981]. It thus suffices to show how $n$ elements in a finite monoid can be multiplied with $\Theta(\log n)$ depth. We show how a log-depth universal transformer can implement the standard binary tree construction [Barrington and Thérien, 1988, Liu et al., 2023, Merrill et al., 2024] where each level multiplies two items, meaning the overall depth is $\Theta(\log\lvert w\rvert)$. We will represent a tree over the input tokens within the transformer. Each level of the tree will take 8 transformer layers. We define a notion of active tokens: at level 0, all tokens are active, and, at level $\ell$, tokens at $t\cdot 2^{\ell}-1$ for any $t$ will remain active, and all other tokens will be marked as inactive. As an invariant, active token $i=t\cdot 2^{\ell}-1$ in level $\ell$ will store a unit-norm vector $\delta^{\ell}_{i}$ that represents the cumulative product of tokens from $i-2^{\ell}+1$ to $i$.
We now proceed by induction over $\ell$, defining the behavior of non-\$ tokens at layers that make up level $\ell$. The current group element $\delta^{\ell}_{i}$ stored at active token $i$ is, by inductive assumption, the cumulative product from $i-2^{\ell}+1$ to $i$. Let $\alpha^{\ell}_{i}$ denote that token $i$ is active. By Lemma 7 we use a layer to store $i-1$ at token $i$. The next layer attends with query $\phi(i-1)$, key $\phi(j)$, and value $\delta^{\ell}_{j}$ to retrieve $\delta^{\ell}_{i-1}$, the group element stored at the previous token. Finally, another layer attends with query $\vec{1}$, key $\langle\phi(j)_{1},\alpha^{\ell}_{i}\rangle$, and value $\delta^{\ell}_{j-1}$ to retrieve the group element $\delta^{\ell}_{j^{*}}$ stored at the previous active token, which represents the cumulative product from $i-2\cdot 2^{\ell}+1$ to $i-2^{\ell}$. Next, we will use two layers to update $\delta^{\ell}_{i}\leftarrow\delta^{\ell+1}_{i}$ and $\delta^{\ell}_{j}\leftarrow\vec{0}$, which is achieved as follows. First, we assert there exists a single feedforward layer that uses a table lookup to compute $\delta^{\ell}_{j^{*}},\delta^{\ell}_{i}\mapsto d$ such that $d/\lVert d\rVert=\delta^{\ell}_{j^{*}}\cdot\delta^{\ell}_{i}=\delta^{\ell+1}_{i}$. Next, we invoke Lemma 6 to construct a layer that adds $d$ to an empty cell of the residual stream and then another layer that deletes it. This second layer can now read both $\delta^{\ell}_{i},\delta^{\ell}_{j^{*}}$ and $\delta^{\ell+1}_{i}$ (from $d$) as input, and we modify it to add $\delta^{\ell+1}_{i}-\delta^{\ell}_{i}$ to $\delta^{\ell}_{i}$, changing its value to $\delta^{\ell+1}_{i}$. Similarly, we modify it to add $-\delta^{\ell}_{j^{*}}$ to $\delta^{\ell}_{j^{*}}$ to set it to 0. A feedforward network then subtracts $\delta^{\ell}_{i}$ from the residual stream and adds $\delta^{\ell}_{i}\cdot\delta^{\ell}_{j}$. This requires at most 4 layers.
To determine activeness in layer $\ell+1$, each token $i$ attends to its left to compute $c_{i}/i$, where $c_{i}$ is the prefix count of active tokens, inclusive of the current token. We then compute $\phi(c_{i}/i,1/i)=\phi(c_{i})$ and store $c_{i}$ it temporarily in the residual stream. At this point, we use Lemma 1 to construct 7 layers that compute $c_{i}\bmod 2$ with no storage overhead. The current token is marked as active in layer $\ell+1$ iff $c_{i}=0\mod 2$, which is equivalent to checking whether $i=t\cdot 2^{\ell}-1$ for some $t$. In addition to updating the new activeness $\alpha^{\ell+1}_{i}$, we also persist store the previous activeness $\alpha^{\ell}_{i}$ in a separate cell of the residual stream and clear $c_{i}$. This requires at most 8 layers.
Appendix D Graph Connectivity Proof
See 2
@@ Appendix D Graph Connectivity Proof
The idea is to use the first $n^{2}$ tokens of the transformer to construct binary predicates $B_{\ell}(i,j)$ for $\ell\in\{0,1,\ldots,\lceil\log n\rceil\}$ capturing whether $G$ has a path of length at most $2^{\ell}$ from $i$ to $j$. To this end, the transformer will use the $n^{3}$ padding tokens to also construct intermediate ternary predicates $C_{\ell}(i,k,j)$ for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$ capturing whether $G$ has paths of length at most $2^{\ell-1}$ from $i$ to $k$ and from $k$ to $j$. These two series of predicates are computed from each other iteratively, as in standard algorithms for graph connectivity:
@@
The crucial part is to construct a transformer that correctly operationalizes the computation of predicates $B_{\ell}$ and $C_{\ell}$. The input to the transformer is the adjacency matrix $A$ represented using $n^{2}$ tokens from $\{0,1\}$, followed by $n^{3}$ padding tokens $\square$, and finally the source and target nodes $s,t\in\{1,\ldots,n\}$ represented in unary notation using special tokens $a$ and $b$:
$A_{1,1}\ldots A_{1,n}\ A_{2,1}\ldots A_{2,n}\ \ldots\ldots\ A_{n,1}\ldots A_{n,n}\ \underbrace{\square\ldots\ldots\ldots\square}_{n^{3}}\ \underbrace{a\ldots\ldots a}_{s}\ \underbrace{b\ldots\ldots b}_{t}$
@@
The transformer uses the next 15 layers to compute and store in the residual stream the semantic “coordinates” of each of the first $n^{2}+n^{3}$ token position as follows. For each of the first $n^{2}$ positions $p=in+j$ with $1\leq p\leq n^{2}$, it uses Lemma 1 (7 layers) with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi(i)$ and $\phi(j)$ to the residual stream at position $p$. In parallel, for each of the next $n^{3}$ positions $p=n^{2}+(in^{2}+kn+j)$ with $n^{2}+1\leq p\leq n^{2}+n^{3}$, it uses Lemma 1 with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi((i+1)n+k)$ and $\phi(j)$ to the residual stream. It then uses the lemma again (7 more layers), this time with $a_{i}$ set to $(i+1)n+k$ and $m$ again set to $n$, to add $\phi(i+1)$ and $\phi(k)$ to the residual stream. Lastly, it uses Lemma 7 applied to $\phi(i+1)$ to add $\phi(i)$ to the residual stream.
Layer 17 of the transformer computes the predicate $B_{0}(i,j)$ at the first $n^{2}$ token positions as follows. At position $p=in+j$, it uses Lemma 8 to compute $\mathbb{I}(\phi(A(i,j)=\phi(1))$ and $\mathbb{I}(\phi(i)=\phi(j))$; note that $\phi(A(i,j))$, $\phi(i)$, and $\phi(j)$ are available in the residual stream at position $p$. It then uses a feedforward layer to output $1$ if both of these are $1$, and output 00 otherwise. This is precisely the intended value of $B_{0}(i,j)$. The sublayer then adds $B_{0}(i,j)$ to the residual stream. The layer also adds to the residual stream the value $1$, which will be used to initialize the boolean that controls layer alternation in the repeated layers as discussed next.
@@
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $(2\ell-1)+8$ of the transformer computes the predicate $C_{\ell}(i,k,j)$ at the set of $n^{3}$ (padding) positions $p=n^{2}+in^{2}+kn+j$, as follows. It uses two heads, one with query $\langle\phi(i),\phi(k)\rangle$ and the other with query $\langle\phi(k),\phi(j)\rangle$. The keys in the first $n^{2}$ positions $q=i^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$, and the values are set to $B_{\ell-1}(i^{\prime},j^{\prime})$. The two heads thus attend solely to positions with coordinates $(i,k)$ and $(k,j)$, respectively, and retrieve boolean values $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$, respectively, stored there in the previous layer. The layer then uses Lemma 8 to compute $\mathbb{I}(B_{\ell-1}(i,k)=1)$ and $\mathbb{I}(B_{\ell-1}(k,j)=1)$, and uses a feedforward layer to output $1$ if both of these checks pass, and output 00 otherwise. This is precisely the intended value of $C_{\ell}(i,k,j)$. If $\ell>1$, the layer replaces the value $C_{\ell-1}(i,k,j)$ stored previously in the residual stream with the new boolean value $C_{\ell}(i,k,j)$ by adding $C_{\ell}(i,k,j)-C_{\ell-1}(i,k,j)$ to the same coordinates of the residual stream. If $\ell=1$, it simply adds $C_{\ell}(i,k,j)$ to the residual stream.